# Electrically Controlled 0-$\pi$ Oscillations and Josephson Giant Magnetoresistor with PT-Symmetric Antiferromagnetic Bilayers

### v2

@@ Abstract
Introduction.—The quest for fault-tolerant quantum computing has spurred intense research into topological superconductors Fu and Kane (2008); Nayak et al. (2008); Qi et al. (2010); Alicea (2012) and phase-controlled Josephson junctions (JJs) Ioffe et al. (1999); Blatter et al. (2001); You et al. (2002); Yamashita et al. (2005); Buzdin (2005); Jørgensen et al. (2007); Gingrich et al. (2016). In conventional JJs, it is known that the supercurrent $I_{s}$ follows $I_{s}=I_{c}\sin(\phi)$, where $\phi$ is the phase difference between superconducting electrodes and $I_{c}$ is the critical current Josephson (1962, 1964). Introducing magnetism into the junction barrier can induce a $\pi$-phase shift, creating a ground-state current-phase relation $I_{s}=I_{c}\sin(\phi+\pi)$—a hallmark of $\pi$-junctions Robinson et al. (2006); Linder et al. (2008). Such junctions are the building blocks of topologically protected superconducting qubits Gladchenko et al. (2009); Kitaev (2006); Gyenis et al. (2021); Paolo et al. (2019); Yamashita et al. (2005); Brooks et al. (2013); Groszkowski et al. (2018); Smith et al. (2020); Guo et al. (2022); Rajpoot et al. (2022).
@@
In this Letter, we first unveil a mechanism of gate-controllable 0-$\pi$ oscillations in lateral JJs using $PT$-symmetric AFM bilayers [see Fig. 1(a)]. When proximitized to a conventional $s$-wave superconductor, the $PT$ symmetry of the system facilitates interlayer spin-singlet pairing via projection onto the low-energy subspace [see Fig. 1(b)]. Crucially, we show that the displacement field in the weak-link region induces an electrically tunable periodic 0-$\pi$ oscillations due to the unique interlayer Cooper pairing. More interestingly, because of the internal Néel order of the spin-layer-locked Cooper pairs, a Josephson giant magnetoresistor can be naturally designed: the supercurrent is strongly suppressed between antiphase interlayer cooper pairs. This occurs because Cooper pair tunneling becomes nearly prohibited when two sides of junctions exhibit antiparallel Néel ordering [see Fig. 1(c)]. Our theoretical analysis provides valuable insights for future studies of unconventional JJs based on $PT$-symmetric AFM bilayers.
Microscopic Model.—To capture the essential low-energy physics of $PT$-symmetric AFM bilayers, we model the system with a minimal Hamiltonian in the fermionic basis $(\psi^{b}_{\uparrow},\psi^{b}_{\downarrow},\psi^{t}_{\uparrow},\psi^{t}_{\downarrow})^{T}$, which reads
@@
A schematic illustration of the designed lateral JJ with gate-defined weak-link is shown in Fig. 1, which has been experimentally realized Rodan-Legrain et al. (2021); de Vries et al. (2021); Diez-Merida et al. (2023) and theoretically explored Xie et al. (2023a); Hu et al. (2023); Xie et al. (2023b). With the gate voltage $V_{TG}$ and $V_{BG}$, the additional Hamiltonian is $\delta H=V(x)\tau_{z}\sigma_{0}$ adding to Eq. (S1). In this case, a layer potential difference $V(x)$ is induced by displacement field with $V(x)=V_{d}\Theta(x)\Theta(L_{0}-x)$. The $PT$ symmetry operator is given by $\hat{S}=\tau_{x}i\sigma_{y}K$ ($K$ is the complex conjugate), while $V(x)$ breaks $PT$ symmetry in the weak-link region and lift the band degeneracy within the weak-link [see Figs. 2(a) and (b)]. In Fig. 1(b) we can see that the Cooper pairs are contained by the $PT$ partners, yielding the interlayer Cooper pairing with the internal Néel order.
Figure 2: (a) In $PT$-symmetric AFM bilayers, an electron on the top layer has its PTPT-partner on the bottom layer, which are degenerate in energy. (b) When $V_{d}\neq 0$, the $PT$ symmetry is broken by lifting the band degeneracy. (c) The numerical and analytical results of the interlayer pairing gap. The numerical result is evaluated from Eq. (2).
@@
Josephson current and 0-$\pi$ oscillations.—The Josephson supercurrent is mediated by an Andreev reflection process involving an incoming spin-up electron from the top layer and an outgoing spin-down hole from the bottom layer. The full junction can be described by the Bogoliubov–de Gennes (BdG) Hamiltonian in the Nambu basis $(\psi,\psi^{\dagger})^{T}$ as
@@
In terms of $\hat{H}(x)$ in Eq. (4), the supercurrent through the junction can be evaluated in the scattering matrix framework by Beenakker (1991); Brouwer and Beenakker (1997); Sun et al. (2024)
@@
where $\omega_{n}=(2n+1)\pi k_{B}T$ are fermionic Matsubara frequencies, $T$ is the temperature. The scattering matrices for the Andreev reflection $\mathcal{S}_{A}$ and the normal scattering $\mathcal{S}_{N}$ can be found in the Supplementary Materials II sup . Here we consider a wide sample $W\gg L_{0}$ with the junction width $W$. So there is the approximate translational invariance along $y$ axis and the preserved wave vector $k_{y}$. Then the Josephson current density can be evaluated by integrating the injection angle $\theta_{i}$, which reads
@@
Josephson giant magnetoresistor.—We now point out another key property of these exotic spin-layer-locked Cooper pairs enabled by $PT$ symmetry: the Josephson giant magnetoresistance (JGMR) effect. In conventional giant magnetoresistor devices, the electrical resistance decreases when adjacent ferromagnetic layers have parallel magnetization alignments and increases significantly for antiparallel configurations [see Fig. 4(a)]. This spin-dependent transport arises from the relative orientation of the ferromagnetic layers’ magnetization, which is the fundamental mechanism behind the tunneling magnetoresistance effect Baibich et al. (1988); Camley and Barnaś (1989); Xiao et al. (1992).
@@
For the detailed derivations, see Appendix A. Then, we can find the JGMR quality factor

### v3

# Unconventional Josephson effects in PT-symmetric antiferromagnetic bilayers

@@ Abstract
I Introduction
The quest for fault-tolerant quantum computing has spurred intense research into topological superconductors Fu and Kane (2008); Nayak et al. (2008); Qi et al. (2010); Alicea (2012) and phase-controlled Josephson junctions (JJs) Ioffe et al. (1999); Blatter et al. (2001); You et al. (2002); Yamashita et al. (2005); Buzdin (2005); Jørgensen et al. (2007); Gingrich et al. (2016). In conventional JJs, it is known that the supercurrent $I_{s}$ follows $I_{s}=I_{c}\sin(\phi)$, where $\phi$ is the phase difference between superconducting electrodes and $I_{c}$ is the critical current Josephson (1962, 1964). Introducing magnetism into the junction barrier can induce a $\pi$-phase shift, creating a ground-state current-phase relation $I_{s}=I_{c}\sin(\phi+\pi)$—a hallmark of $\pi$-junctions Robinson et al. (2006); Linder et al. (2008). Such junctions are the building blocks of topologically protected superconducting qubits Gladchenko et al. (2009); Kitaev (2006); Gyenis et al. (2021); Paolo et al. (2019); Yamashita et al. (2005); Brooks et al. (2013); Groszkowski et al. (2018); Smith et al. (2020); Guo et al. (2022); Rajpoot et al. (2022).
@@ I Introduction
In this work, we first unveil a mechanism of gate-controllable 0-$\pi$ oscillations in lateral JJs using $PT$-symmetric AFM bilayers [see Fig. 1(a)]. When proximitized to a conventional $s$-wave superconductor, the $PT$ symmetry of the system facilitates interlayer spin-singlet pairing via projection onto the low-energy subspace [see Fig. 1(b)]. Crucially, we show that the displacement field in the weak-link region induces an electrically tunable periodic 0-$\pi$ oscillations due to the unique interlayer Cooper pairing. More interestingly, because of the internal Néel order of the spin-layer-locked Cooper pairs, a Josephson giant magnetoresistor can be naturally designed: the supercurrent is strongly suppressed between antiphase interlayer cooper pairs. This occurs because Cooper pair tunneling becomes nearly prohibited when two sides of junctions exhibit antiparallel Néel ordering [see Fig. 1(c)]. Our theoretical analysis provides valuable insights for future studies of unconventional JJs based on $PT$-symmetric AFM bilayers.
II Microscopic Model
To capture the essential low-energy physics of $PT$-symmetric AFM bilayers, we model the system with a minimal Hamiltonian in the fermionic basis $(\psi^{b}_{\uparrow},\psi^{b}_{\downarrow},\psi^{t}_{\uparrow},\psi^{t}_{\downarrow})^{T}$, which reads
@@ II Microscopic Model
A schematic illustration of the designed lateral JJ with gate-defined weak-link is shown in Fig. 1, which has been experimentally realized Rodan-Legrain et al. (2021); de Vries et al. (2021); Diez-Merida et al. (2023) and theoretically explored Xie et al. (2023a); Hu et al. (2023); Xie et al. (2023b). With the gate voltage $V_{TG}$ and $V_{BG}$, the additional Hamiltonian is $\delta H=V(x)\tau_{z}\sigma_{0}$ adding to Eq. (1). In this case, a layer potential difference $V(x)$ is induced by displacement field with $V(x)=V_{d}\Theta(x)\Theta(L_{0}-x)$. The $PT$ symmetry operator is given by $\hat{S}=\tau_{x}i\sigma_{y}K$ ($K$ is the complex conjugate), while $V(x)$ breaks $PT$ symmetry in the weak-link region and lift the band degeneracy within the weak-link [see Figs. 2(a) and (b)]. In Fig. 1(b) we can see that the Cooper pairs are contained by the $PT$ partners, yielding the interlayer Cooper pairing with the internal Néel order.
Figure 2: (a) In $PT$-symmetric AFM bilayers, an electron on the top layer has its PTPT-partner on the bottom layer, which are degenerate in energy. (b) When $V_{d}\neq 0$, the $PT$ symmetry is broken by lifting the band degeneracy. (c) The numerical and analytical results of the interlayer pairing gap. The numerical result is evaluated from Eq. (8). (d) Schematic picture of the Andreev reflections of lateral JJ built by AFM bilayers. The Andreev reflection process involves an incoming spin-up electron from the top layer and an outgoing spin-down hole from the bottom layer.
@@
In $PT$-symmetric AFM bilayers, the wavefunctions for each layer without the interlayer coupling $g$ can be written as
$\psi_{t\uparrow}$ $=(0,0,e^{-i\frac{\varphi}{2}}\cos\frac{\theta}{2},e^{i\frac{\varphi}{2}}\sin\frac{\theta}{2})^{T}$ (2)
$\psi_{t\downarrow}$ $=(0,0,-e^{i\frac{\varphi}{2}}\sin\frac{\theta}{2},e^{i\frac{\varphi}{2}}\cos\frac{\theta}{2})^{T}$ (3)
and
$\psi_{b\uparrow}$ $=(e^{-i\frac{\varphi}{2}}\cos\frac{\theta}{2},e^{i\frac{\varphi}{2}}\sin\frac{\theta}{2},0,0)^{T}$ (4)
$\psi_{b\downarrow}$ $=(-e^{-i\frac{\varphi}{2}}\sin\frac{\theta}{2},e^{i\frac{\varphi}{2}}\cos\frac{\theta}{2},0,0)^{T}.$ (5)
Here $b(t)$ denotes the bottom (top) layer. Focusing on the low-energy states localized on the top layer-spin up and bottom layer-spin down, thus in the presence of $g$ the wavefunctions are perturbed to
$u_{1}=\frac{1}{N_{0}}(\psi_{t\uparrow}-\frac{g}{2J_{ex}}\psi_{b\uparrow})$ (6)
$u_{2}=\frac{1}{N_{0}}(\psi_{b\downarrow}-\frac{g}{2J_{ex}}\psi_{t\downarrow}),$ (7)
where $N_{0}=\sqrt{1+g^{2}/4J_{ex}^{2}}$ is the normalization factor. We can obtain the transformation matrix in the reduced Hilbert space $\tilde{U}(\theta,\varphi)=[u_{1}(\theta,\varphi),u_{2}(\theta,\varphi)]$.
@@
III Josephson current and 0-$\pi$ oscillations.
The Josephson supercurrent is mediated by an Andreev reflection process involving an incoming spin-up electron from the top layer and an outgoing spin-down hole from the bottom layer. The full junction can be described by the Bogoliubov–de Gennes (BdG) Hamiltonian in the Nambu basis $(\psi,\psi^{\dagger})^{T}$ as
@@
As shown in Fig. 2(d), the incoming wavefunction $\psi_{\text{in }}=[c_{e}^{-}(L),c_{h}^{+}(L),c_{e}^{+}(R),c_{h}^{-}(R)]^{T}$ and the outcoming wavefunction $\psi_{\text{out }}=[c_{e}^{+}(L),c_{h}^{-}(L),c_{e}^{-}(R),c_{h}^{+}(R)]^{T}$ satisfy the relations $\psi_{\text{out }}=\mathcal{S}_{\mathrm{A}}\psi_{\text{in }}$ and $\psi_{\text{in }}=\mathcal{S}_{\mathrm{N}}\psi_{\text{out }}$. The scattering matrices of the junction for the Andreev reflection $\mathcal{S}_{A}$ and the normal scattering $\mathcal{S}_{N}$ read
$\mathcal{S}_{\mathrm{A}}=\left(\begin{array}[]{cccc}0&e^{i\frac{\phi}{2}-i\chi}&0&0\\ e^{-i\frac{\phi}{2}-i\chi}&0&0&0\\ 0&0&0&e^{-i\frac{\phi}{2}-i\chi}\\ 0&0&e^{i\frac{\phi}{2}-i\chi}&0\end{array}\right),$ (11)
and
$\mathcal{S}_{\mathrm{N}}=\left(\begin{array}[]{cccc}0&0&e^{-ik_{e,l-}L_{0}}&0\\ 0&0&0&e^{-ik_{h,l+}L_{0}}\\ e^{ik_{e,l+}L_{0}}&0&0&0\\ 0&e^{ik_{h,l-}L_{0}}&0&0\end{array}\right)$ (12)
Here we take $L_{0}$ the junction length. If we only consider the linear order correction of Fermi momenta induced by the displacement field $V_{d}$, the correction can be written as $k_{e,l\alpha}\approx k_{N,l\alpha}^{0}+\delta k_{e,l\alpha},k_{h,l\alpha}\approx k_{N,l\alpha}^{0}+\delta k_{h,l\alpha}$ in which $k_{N,l\alpha}^{0}=\sqrt{\mu/\lambda}\cos\theta_{i}$ and
$\delta k_{e,l\alpha}=\frac{\epsilon-\delta\varepsilon_{l}}{v_{f}^{\alpha}},\\ \\ \delta k_{h,l\alpha}=\frac{\epsilon-\delta\varepsilon_{l}}{-v_{f}^{\alpha}}$ (13)
Here the energy shift $\delta\varepsilon_{l}=l(V_{d}+g_{s}\mu_{B}\bm{n}\cdot\bm{B})$ contains both the displacement field and magnetic field. Here $g_{s}$ is the Landé g-factor. In terms of $\hat{H}(x)$ in Eq. (10), the supercurrent through the junction can be evaluated in the scattering matrix framework by Beenakker (1991); Brouwer and Beenakker (1997); Sun et al. (2024)
@@
where $\omega_{n}=(2n+1)\pi k_{B}T$ are fermionic Matsubara frequencies, $T$ is the temperature. Here we consider a wide sample $W\gg L_{0}$ with the junction width $W$. So there is the approximate translational invariance along $y$ axis and the preserved wave vector $k_{y}$. Then the Josephson current density can be evaluated by integrating the injection angle $\theta_{i}$, which reads
@@
Figure 5: Crystal structure and band structure of VSe2. (a) The crystal structure of bilayer VSe2. The red and green spheres represent V and Se atoms, respectively. (b) The band structure of bilayer VSe2 with projections on the $d_{xy}$ and $d_{z^{2}}$ orbitals of V atoms in different layers. The size of the markers represents the weight of the corresponding orbitals.
@@
IV Josephson giant magnetoresistor
We now point out another key property of these exotic spin-layer-locked Cooper pairs enabled by $PT$ symmetry: the Josephson giant magnetoresistance (JGMR) effect. In conventional giant magnetoresistor devices, the electrical resistance decreases when adjacent ferromagnetic layers have parallel magnetization alignments and increases significantly for antiparallel configurations [see Fig. 4(a)]. This spin-dependent transport arises from the relative orientation of the ferromagnetic layers’ magnetization, which is the fundamental mechanism behind the tunneling magnetoresistance effect Baibich et al. (1988); Camley and Barnaś (1989); Xiao et al. (1992).
@@ IV Josephson giant magnetoresistor
For the detailed derivations, see Supplementary Materials. Then, we can find the JGMR quality factor